# METDrive: Multi-modal End-to-end Autonomous Driving with Temporal Guidance

### v1

@@ V Conclusion and Future Work
Through this paper, we identify a common shortcoming in existing end-to-end systems for the CARLA simulator: the absence of guidance from additional modalities, which leads to failure cases in long-term tasks. To address this issue, we propose utilizing ego-related temporal cues, such as rotation angles, steering, throttle signals, and waypoint vectors, to guide the geometric features derived from perception sensors in predicting waypoints. Consequently, we design a temporal guidance loss to monitor this integration process. To evaluate our system, we conducted experiments on the CARLA leaderboard’s Longest6 benchmark, achieving $70\%$ driving score, $94\%$ route completion score, and $0.78$ infraction score.
Despite employing time series features as temporal guidance within this study, the efficacy of the feature fusion approach is constrained by the quality of the ego state inputs derived from the collected dataset. Specifically, given that the dataset was generated using an autonomous agent within the CARLA simulation environment, the agent’s performance was characterized by inconsistencies, resulting in a lack of smoothness and natural driving behavior. Consequently, the ego state data were replete with significant noise, preprocessing steps such as filtering before they could be utilized effectively for our training.
In light of these limitations, our future research routine will focus on the acquisition of a higher fidelity dataset. This dataset will be sourced from expert human drivers navigating a diverse range of traffic scenarios. By incorporating data from skilled human operators, we anticipate a reduction in the aforementioned noise and a resultant enhancement in the quality of the input data. Additionally, we intend to develop a more robust framework that encompasses a broader spectrum of time series encoders and feature fusion tailored for ego state inputs. This enhanced approach is expected to mitigate the current issues and improve the overall performance and reliability of the system under development.

### v2

@@ V Conclusion and Future Work
In this paper, we identify a common shortcoming in existing end-to-end systems for the CARLA simulator: the absence of guidance from additional modalities, which may lead to failure cases in long-term tasks. To address this issue, we propose to utilize ego-related temporal cues, such as rotation angles, steering, throttle signals, and waypoint vectors, to guide the geometric features derived from perception sensors in predicting waypoints. Consequently, we design a temporal guidance loss to monitor this integration process. To evaluate our system, we conducted experiments on the CARLA Longest6 benchmark, achieving a $70\%$ driving score, a $94\%$ route completion score and a $0.78$ infraction score.
Despite time series features are used as temporal guidance in this study, the efficacy of the feature fusion approach is constrained by the quality of the ego state inputs derived from the collected dataset. Specifically, given that the dataset was generated using an autonomous agent within the CARLA simulation environment, the agent’s performance was characterized by inconsistencies, resulting in a lack of smoothness and natural driving behavior. Consequently, the ego state data were replete with significant noise, preprocessing steps such as filtering before they could be utilized effectively for our training.
In light of these limitations, our future research routine will focus on the acquisition of a higher fidelity dataset. This dataset will be sourced from expert human drivers who navigate a diverse range of traffic scenarios. By incorporating data from skilled human operators, we anticipate a reduction in the aforementioned noise and a resultant enhancement in the quality of the input data. Additionally, we intend to develop a more robust framework that encompasses a broader spectrum of time series encoders and feature fusion tailored for ego state inputs. This enhanced approach is expected to mitigate current issues and improve the overall performance and reliability of the system under development.